OpenStudy (anonymous):

Solve. 8x = 4x – 32 A. 4 B. 8 C. –4 D. –8

Mehek (mehek14):

subtract 8x - 4x

OpenStudy (anonymous):

so C

Mehek (mehek14):

no

OpenStudy (midhun.madhu1987):

Subtract 4x from both sides. Divide by 4 on both sides. You get the answer..

Mehek (mehek14):

first tell me what 8x - 4x =

OpenStudy (anonymous):

no, after you get 4x divide 4 by -32 and that is your answer.

OpenStudy (anonymous):

4

Mehek (mehek14):

4?

OpenStudy (anonymous):

no its d

OpenStudy (anonymous):

to your question

Mehek (mehek14):

4x not 4

Mehek (mehek14):

8x - 4x = 4x

OpenStudy (anonymous):

-32/4 = -8

OpenStudy (freckles):

\[8x=4x-32 \\ \text{ subtract } 4x \text{ on both sides } \\ 8x-4x=-32 \\ 4x=-32 \\ \text{ last step is to divide } 4 \text{ on both sides }\]

Mehek (mehek14):

\(4x = -32\) divide \(\dfrac{-32}{4}\)

OpenStudy (anonymous):

exactly spear

OpenStudy (anonymous):

thanks for your help guys

OpenStudy (anonymous):

welcome

OpenStudy (skullpatrol):

8x = 4x - 32 8x - 4x = -32 4x = -32 x = -8 Check: 8(-8) =?= 4(-8) - 32 -64 =?= -32 - 32 -64 = -64 {-8} Answer